# Cosmological dynamics of interacting dark energy and dark matter in $f(Q)$ gravity

### v1

@@ IV.2.2 $R_{2}$: Non-metricity dominated fixed point
The characteristics of this critical point solely depend on the model parameter $n$. Corresponding to this fixed point, the matter density, dark energy density, and effective EoS parameters are obtained as
@@
The corresponding eigenvalues for the fixed point $R_{2}$ are as follows:
{4,  $\{{4,\,\,1-\frac{\alpha}{2n}\}}.$ (32)
We determine the stability of this fixed point for various ranges of $\alpha$. The ranges are as follows:
@@ IV.2.3 $R_{3}$: Scaling solution point
Corresponding to this fixed point, the matter density, dark energy density, and effective EoS parameters read as
@@ IV.2.4 $R_{4}$: Scaling solution point
The fixed point $R_{4}$ represents the scaling solution point of the universe. In addition, it is worth noting that our result modifies the scaling solution with the interacting dark energy parameter $\alpha$. The point $R_{4}$ is
@@
Corresponding to this fixed point, the matter density, dark energy density, and effective EoS parameters are obtained as
@@
One can note that all three parameters depend on the model parameter $n$ as well as the interacting parameter $\alpha$. Check appendix A for calculations. The conditions for an accelerating universe and the stability are shown in Figs. 6 and 7.
Figure 6: This plot shows the relationship between the model parameter $n$ and the interacting parameter $\alpha$ for the accelerating universe. The light blue region represents the quintessence-like behavior of an accelerating universe (i.e., $-1<\omega<-1/3$), while the golden region represents the phantom-like behavior of an accelerating universe (i.e., $\omega<-1$).
Figure 7: The region plot illustrates the stable, unstable, and saddle behaviors associated with the fixed point $R_{4}$.
@@ IV.2.5 $R_{5}$: Additional scaling solution point
the matter density, dark energy density, and effective EoS parameters are given by
@@
The conditions for an accelerating universe and the stability are depicted in Figs. 8 and 9.
Figure 8: The plot show the relationship between the model parameter $n$ and the interacting parameter $\alpha$ for the accelerating universe. The light blue region represents the quintessence-like behavior of an accelerating universe (i.e., $-1<\omega<-1/3$), while the golden region represents the phantom-like behavior of an accelerating universe (i.e., $\omega<-1$).
Figure 9: The region plot illustrates the stable, unstable, and saddle behaviors associated with the fixed point $R_{5}$.
In Figures 10 and 11, we illustrate the evolution of $\Omega_{m}$, $\Omega_{r}$, $\Omega_{DE}$, $q$, and $\omega$ for different values of the model parameter $n$ and the interaction parameter $\alpha$. For $n=3/2$ and $\alpha=4$ in Figure 10, $\alpha>0$ indicates that the coupling term $\mathcal{U}>0$, signifying energy transfer from dark matter to dark energy. This figure shows that the universe was dominated by matter in the early stages and will be dominated by dark energy in later stages. Currently, the universe is dominated by dark energy, with parameter values $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, and $\Omega_{DE}=0.7$, $q_{0}=-0.548$, and $\omega_{0}=-0.696$. For these values, the fixed points $R_{1}$ and $R_{4}$ are stable and represent the de Sitter and quintessence acceleration solutions, respectively.
In Figure 11, with $n=3/2$ and $\alpha=-4$, $\alpha<0$ indicates that the coupling term $\mathcal{U}<0$, meaning energy transfers from dark energy to dark matter. This figure shows that the universe was dominated by dark energy in the early stages and will be dominated by dark matter in later stages. Currently, the universe remains dominated by dark energy, with parameter values $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, and $\Omega_{DE}=0.7$, $q_{0}=-0.55$, and $\omega_{0}=-0.70$. For these values, the fixed points $R_{4}$ and $R_{5}$ have imaginary values, and the remaining point cannot exhibit stable behavior.
Figure 10: The upper panel displays density parameters for matter ($\Omega_{m}$), dark energy ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift $z$. The lower panel shows the deceleration parameter $q$ and the equation of state parameter $\omega$ as functions of redshift $z$. These two panels illustrate the trajectory for the positive coupling parameter, which can transfer energy from dark matter to dark energy.
Figure 11: The left panel displays density parameters for matter ($\Omega_{m}$), dark energy ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift $z$. The right panel shows the deceleration parameter $q$ and the equation of state parameter $\omega$ as functions of redshift $z$. These two panels illustrate the trajectory for the negative coupling parameter, which can transfer energy from dark energy to dark matter.
Figure 12: Phase plot of $x$ vs $y$. The phase plot above shows the stable behavior of $R_{1}$ and $R_{4}$ for $\alpha>0$.
@@ V Conclusion
In this work, we have explored the behavior of interacting dark energy and dark matter within $f(Q)$ gravity, employing a standard framework of dynamical system analysis. We have considered the power-law $f(Q)=6\gamma\,H_{0}^{2}\left(\tfrac{Q}{Q_{0}}\right)^{n}$ model incorporating with two different forms of interacting dark energy and dark matter: $3\alpha H\rho_{m}$ and $\frac{\alpha}{3H}\rho_{m}\rho_{DE}$. The parameter $\alpha$ in the interacting terms plays a crucial role in determining the viable conditions and estimating the transition from the matter-dominated era (saddle point) to the dark energy-dominated era (stable node) in viable gravity models at late times. As a result, we have discovered fixed points that can be represented as the late-time accelerating universe in $f(Q)$ gravity. For the form of ${\cal U}=3\alpha H\rho_{m}$, we have illustrated the evolution of $\Omega_{m}$, $\Omega_{r}$, $\Omega_{DE}$, $q$, and $\omega$ for different values of the model parameter $n$ and the interaction parameter $\alpha$. For $n=3/2$ and $\alpha=1/2$ and $\alpha>0$, we found that the coupling term $\mathcal{U}>0$, signifying energy transfer from dark matter to dark energy implying that the universe was dominated by matter in the early stages and would be dominated by dark energy in later stages. With the current data, fixed point $P_{1}$ is stable and represents the de Sitter acceleration solution, while fixed point $P_{2}$ is a saddle node, and $P_{3}$ is an unstable node that cannot demonstrate universal acceleration. Moreover, we have considered another situation of which $n=3/2$ and $\alpha=-2$, $\alpha<0$. We discovered for the coupling term $\mathcal{U}<0$ that energy transfers from dark energy to dark matter. In this case, the universe was dominated by dark energy in the early stages and will be dominated by dark matter in later stages. With the current data, fixed point $P_{3}$ is stable and exhibits acceleration for the early and present universe but fails to show acceleration for late times, while fixed point $P_{1}$ is a saddle node, and $P_{2}$ is an unstable node.
@@
For the form of ${\cal U}=\frac{\alpha}{3H}\rho_{m}\rho_{DE}$, the evolution of $\Omega_{m}$, $\Omega_{r}$, $\Omega_{DE}$, $q$, and $\omega$ for different values of the model parameter $n$ and the interaction parameter $\alpha$ have been examined. We have considered $n=3/2$ and $\alpha=4$ ($\alpha>0$) and found that the coupling term $\mathcal{U}>0$, signifying energy transfer from dark matter to dark energy. Our results show that the universe was dominated by matter in the early stages and would be dominated by dark energy in later stages. Using the observational data, the fixed points $R_{1}$ and $R_{4}$ were stable and would represent the de Sitter and quintessence acceleration solutions. Additionally, for $n=3/2$ and $\alpha=-4$ ($\alpha<0$), we found that the coupling term $\mathcal{U}<0$, meaning energy transfers from dark energy to dark matter implying that the universe was dominated by dark energy in the early stages and would be dominated by dark matter in later stages. Employing the current data, the fixed points $R_{4}$ and $R_{5}$ have imaginary values, and the remaining point cannot exhibit stable behavior.
@@
Although many viable $f(Q)$ models explain the dark energy problem in cosmology, our qualitative results from this work can serve as guidelines for more detailed studies. They can also be used as complementary constraints on viable $f(Q)$ models, alongside other cosmological constraints on $f(Q)$ theories. Our framework, based on the cosmological dynamics of interacting dark energy and dark matter in $f(Q)$ gravity, constitutes the natural template for beyond the standard gravity model, for example, Teleparallel gravity and Gauss-Bonnet gravity, or even more generalizations.
@@ Acknowledgments
GNG acknowledges University Grants Commission (UGC), New Delhi, India for awarding a Senior Research Fellowship (UGC-Ref. No.: 201610122060). PKS acknowledges Science and Engineering Research Board, Department of Science and Technology, Government of India for financial support to carry out Research project No.: CRG/2022/001847 and IUCAA, Pune, India for providing support through the visiting Associateship program.

### v2

@@ IV.2.2 $R_{2}$: Non-metricity dominated fixed point
The characteristics of this critical point solely depend on the model parameter $n$. Corresponding to this fixed point, the matter density, DE density, and effective EoS parameters are obtained as
@@
The corresponding eigenvalues for the fixed point $R_{2}$ are {4,  $\{{4,\,\,1-\frac{\alpha}{2n}\}}$. We determine the stability of this fixed point for various ranges of $\alpha$. The ranges are as follows
@@ IV.2.3 $R_{3}$: Scaling solution point
Corresponding to this fixed point, the matter density, DE density, and effective EoS parameters read as
@@ IV.2.4 $R_{4}$: Scaling solution point
The fixed point $R_{4}$ represents the scaling solution point of the universe. In addition, it is worth noting that our result modifies the scaling solution with the interacting DE parameter $\alpha$. The point $R_{4}$ is
@@
Corresponding to this fixed point, the matter density, DE density, and effective EoS parameters are obtained as
@@
One can note that all three parameters depend on the model parameter $n$ as well as the interacting parameter $\alpha$. Check appendix A for calculations. The conditions for an accelerating universe and the stability are shown in Figs. 6(a) and 6(b).
Figure 6: The region plot (a) shows the relationship between the model parameter $n$ and the interacting parameter $\alpha$ for the accelerating universe. The light blue region represents the quintessence-like behavior of an accelerating universe (i.e., $-1<\omega<-1/3$), while the golden region represents the phantom-like behavior of an accelerating universe (i.e., $\omega<-1$). The region plot (b) illustrates the stable, unstable, and saddle behaviors associated with the fixed point $R_{4}$.
@@ IV.2.5 $R_{5}$: Additional scaling solution point
the matter density, DE density, and effective EoS parameters are given by
@@
The conditions for an accelerating universe and the stability are depicted in Figs. 7(a) and 7(b).
Figure 7: The region plot (a) shows the relationship between the model parameter $n$ and the interacting parameter $\alpha$ for the accelerating universe. The light blue region represents the quintessence-like behavior of an accelerating universe (i.e., $-1<\omega<-1/3$), while the golden region represents the phantom-like behavior of an accelerating universe (i.e., $\omega<-1$). The region plot (b) illustrates the stable, unstable, and saddle behaviors associated with the fixed point $R_{5}$.
In Figures 8 and 9, we illustrate the evolution of $\Omega_{m}$, $\Omega_{r}$, $\Omega_{DE}$, $q$, and $\omega$ for different values of the model parameter $n$ and the interaction parameter $\alpha$. For $n=3/2$ and $\alpha=4$ in Figure 8, $\alpha>0$ indicates that the coupling term $\mathcal{U}>0$, signifying energy transfer from DM to DE. This figure shows that the universe was dominated by matter in the early stages and will be dominated by DE in later stages. Currently, the universe is dominated by DE, with parameters value $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, $\Omega_{DE}=0.7$, $q_{0}=-0.548$, and $\omega_{0}=-0.696$. For these values, the fixed points $R_{1}$ and $R_{4}$ are stable and represent the de Sitter and quintessence acceleration solutions, respectively.
In Figure 9, with $n=3/2$ and $\alpha=-4$, $\alpha<0$ indicates that the coupling term $\mathcal{U}<0$, meaning energy transfers from DE to DM. This figure shows that the universe was dominated by DE in the early stages and will be dominated by DM in later stages. Currently, the universe remains dominated by DE, with parameters value $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, $\Omega_{DE}=0.7$, $q_{0}=-0.55$, and $\omega_{0}=-0.70$. For these values, the fixed points $R_{4}$ and $R_{5}$ have imaginary values, and the remaining point cannot exhibit stable behavior.
Figure 8: The left panel displays density parameters for matter ($\Omega_{m}$), DE ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift $z$. The right panel shows the deceleration parameter $q$ and the EoS parameter $\omega$ as functions of redshift $z$. These two panels illustrate the trajectory for the positive coupling parameter, which can transfer energy from DM to DE.
Figure 9: The left panel displays density parameters for matter ($\Omega_{m}$), DE ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift $z$. The right panel shows the deceleration parameter $q$ and the EoS parameter $\omega$ as functions of redshift $z$. These two panels illustrate the trajectory for the negative coupling parameter, which can transfer energy from DE to DM.
Figure 10: Phase plot of $x$ versus $y$. These phase plots depict the stable behavior of $R_{1}$ and $R_{4}$ for $\alpha>0$.
@@ V Conclusion
In this work, we have explored the behavior of interacting DE and DM within $f(Q)$ gravity, employing a standard framework of dynamical system analysis. We have considered the power-law $f(Q)=6\gamma\,H_{0}^{2}\left(\tfrac{Q}{Q_{0}}\right)^{n}$ model incorporating with two different forms of interacting DE and DM: $3\alpha H\rho_{m}$ and $\frac{\alpha}{3H}\rho_{m}\rho_{DE}$. The parameter $\alpha$ in the interacting terms plays a crucial role in determining the viable conditions and estimating the transition from the matter-dominated era (saddle point) to the DE-dominated era (stable node) in viable gravity models at late times. As a result, we have discovered fixed points that can be represented as the late-time accelerating universe in $f(Q)$ gravity. For the form of ${\cal U}=3\alpha H\rho_{m}$, we have illustrated the evolution of $\Omega_{m}$, $\Omega_{r}$, $\Omega_{DE}$, $q$, and $\omega$ for different values of the model parameter $n$ and the interaction parameter $\alpha$. For $n=3/2$ and $\alpha=1/2$ and $\alpha>0$, we found that the coupling term $\mathcal{U}>0$, signifying energy transfer from DM to DE implying that the universe was dominated by matter in the early stages and would be dominated by DE in later stages. With the current data, fixed point $P_{1}$ is stable and represents the de Sitter acceleration solution, while fixed point $P_{2}$ is a saddle node, and $P_{3}$ is an unstable node that cannot demonstrate universal acceleration. Moreover, we have considered another situation of which $n=3/2$ and $\alpha=-2$, $\alpha<0$. We discovered for the coupling term $\mathcal{U}<0$ that energy transfers from DE to DM. In this case, the universe was dominated by DE in the early stages and will be dominated by DM in later stages. With the current data, fixed point $P_{3}$ is stable and exhibits acceleration for the early and present universe but fails to show acceleration for late times, while fixed point $P_{1}$ is a saddle node, and $P_{2}$ is an unstable node.
@@
For the form of ${\cal U}=\frac{\alpha}{3H}\rho_{m}\rho_{DE}$, the evolution of $\Omega_{m}$, $\Omega_{r}$, $\Omega_{DE}$, $q$, and $\omega$ for different values of the model parameter $n$ and the interaction parameter $\alpha$ have been examined. We have considered $n=3/2$ and $\alpha=4$ ($\alpha>0$) and found that the coupling term $\mathcal{U}>0$, signifying energy transfer from DM to DE. Our results show that the universe was dominated by matter in the early stages and would be dominated by DE in later stages. Using the observational data, the fixed points $R_{1}$ and $R_{4}$ were stable and would represent the de Sitter and quintessence acceleration solutions. Additionally, for $n=3/2$ and $\alpha=-4$ ($\alpha<0$), we found that the coupling term $\mathcal{U}<0$, meaning energy transfers from DE to DM implying that the universe was dominated by DE in the early stages and would be dominated by DM in later stages. Employing the current data, the fixed points $R_{4}$ and $R_{5}$ have imaginary values, and the remaining point cannot exhibit stable behavior.
@@
Although many viable $f(Q)$ models explain the DE problem in cosmology, our qualitative results from this work can serve as guidelines for more detailed studies. They can also be complementary constraints on viable $f(Q)$ models alongside other cosmological constraints on $f(Q)$ theories. Our framework, based on the cosmological dynamics of interacting DE and DM in $f(Q)$ gravity, constitutes the natural template beyond the standard gravity model, for example, Teleparallel gravity and Gauss-Bonnet gravity, or even more generalizations. Advancing our understanding requires developing new theoretical tools to analyze how interactions influence linear and non-linear regimes. Precisely determining an interaction kernel is crucial, as it could offer profound insights into fundamental physics, particularly the nature and properties of DM and DE.
@@ Acknowledgments
GNG acknowledges University Grants Commission (UGC), New Delhi, India for awarding a Senior Research Fellowship (UGC-Ref. No.: 201610122060). SA acknowledges the Japan Society for the Promotion of Science (JSPS) for providing a postdoctoral fellowship during 2024-2026 (JSPS ID No.: P24318). This work of SA is supported by the JSPS KAKENHI Grant (Number: 24KF0229). PKS acknowledges Anusandhan National Research Foundation, Department of Science and Technology, Government of India for financial support to carry out Research project No.: CRG/2022/001847 and IUCAA, Pune, India for providing support through the visiting Associateship program. We are very much grateful to the honorable referee and to the editor for the illuminating suggestions that have significantly improved our work in terms of research quality, and presentation.